# Eliciting and Distinguishing Between Weak and Incomplete Preferences: Theory, Experiment and Computation

### v5

@@ 4.1         Choice Sizes
We start with a key new variable of interest, which we refer to as the subjects’ choice size, and which is defined simply as the number of gift-card pairs that were chosen at each menu. This variable ranges between 0 and 4 in the Non-Forced-Choice treatment, and between 1 and 4 in the Forced-Choice treatment. However, because choice sizes 1 and 2 (as well as 0 in the former treatment) are always feasible, whereas 3 and 4 are not, we adjust their relative frequencies accordingly. Specifically, Figure 1 presents the distributions of menu-size adjusted choice sizes in the two treatments, which are derived once their absolute frequencies are divided by the total number of menus where these choice sizes might be observed. For example, the denominators here are $50\times N$ and $15\times N$ for choice sizes 1 and 4, respectively, where $N$ is the total number of subjects in the relevant treatment.
This comparison shows that the modal adjusted choice size was 1 in both treatments, at a rate of just over 50%, and was followed by 2 and 3. The deferral rate in the Non-Forced-Choice treatment, defined as the relative frequency of a zero choice size, is 7.4%. This is slightly higher than the 6.9% choose-everything rate, which is obtained by weighted-averaging the relative frequencies where $n$ out of $n$ items were chosen, for $n=2,3,4$ (Table 4). The latter rate is also lower than the corresponding one in the Forced-Choice treatment (8.5%), and the difference is statistically significant ($p<0.001$; two-sided Fisher’s exact test). This difference is consistent with the intuition that, in the absence of the possibility to delay choice when faced with a potentially difficult decision, subjects are more likely to choose everything that is available in the menu, either as a result of following a more general decision rule or, in the context of our experimental design, possibly due to a preference for randomization. Finally, the rate at which subjects chose all 4 alternatives was very low in both treatments, at 3.2% and 2.5% in the Forced- and Free-Choice treatments, respectively.
@@ 4.2         Response Times
Notably, the average response time in both treatments was always lowest ($\approx$ 6 secs) when subjects chose a single pair of gift cards, while it increased monotonically (and statistically significantly) in similar ways as the size of choices and menus increased (Figure 2). These novel empirical facts might be seen as intuitive evidence suggesting that the subjects’ decision was easier at menus where they chose a single gift-card pair, perhaps because that was their clearly preferred one. The latter explanation would be in line with the one often mentioned in the literature of binary forced-choice decisions via sequential-sampling processes such as the drift-diffusion model and extensions (Ratcliff and McKoon, 2008; Alós-Ferrer et al., 2021), although the novelty of identifying a possibly structure for the evolution of this process in free- and non-binary choices would remain.
@@ 4.3         Choice Proportions
Additionally, 69 subjects (51.1%) in the Free-Choice treatment deferred at least once, with deferrals per subject ranging between 1 and 42 (Figure 3-(b); bottom panel), and with a mean/median occurrence of 7.28/5 menus. Moreover, 81 subjects in this treatment (60%) chose all feasible gift-card pairs at least once, with such occurrences per subject ranging between 1 and 38 menus (Figure 3-(b); middle panel), and with a mean/median occurrence of 3.42/1 menus. By contrast, in the Forced-Choice treatment there were 98 subjects (71%) who chose everything at least once, with occurrences per subject ranging between 1 and 45 (Figure 3-(a); middle panel), and with a mean/median of 4.22/2 menus. The difference in the two proportions is borderline (in)significant at the 5% level ($p=0.058$; two-sided Fisher’s exact test). These findings lend further support to the possibility that, unable to delay their choice when faced with a hard decision, subjects are more likely to choose everything.
@@ 4.4         Choice Consistency
For this analysis –which is summarized in Table 5– we compute in [Uncaptioned image]and then compare the Jaccard-Houtman-Maks (JHM) scores of subjects’ active choices while accounting for the possibility of non-trivial indifferences. That is, we compare subjects’ single- and/or multi-valued choices –ignoring any deferral decisions– that are compatible with indifference-permitting utility maximization. First, we focus on the distributions of the JHM scores across the two treatments. Their mean (5.38 v 4.43), median (4.96 v 3.50) and standard deviation (4.02 v 3.75) are uniformly and significantly ($p=0.045$) lower in the Free-Choice than in the Forced-Choice treatment. Thus, the number of partially or fully “mistaken” –from the point of view of utility maximization– active choices is significantly smaller for subjects who were not asked to always choose some gift-card pair(s). This is also reflected in our second test for treatment effects in active-choice consistency, where we compare the proportions of subjects who made perfectly (JHM score of 0) or approximately (JHM score less than or equal to 1, 2, 3, 4 or 5) consistent active choices in the two treatments. There were approximately 4% and 13% perfectly consistent subjects in the Forced- and Free-Choice treatments, respectively ($p=0.007$). The pattern is similar for approximately consistent subjects, with 14.5% v 26% making up to one fully mistaken choice ($p=0.023$), 27.5% v 37% up to two ($p=0.092$), 36% v 50% up to three ($p=0.028$), 43% v 54% up to four ($p=0.069$) and 51% v 59% up to five ($p=0.181$).
@@
Forced Choice Free Choice $p$-value Mean JHM 5.38 4.43 0.045 Median JHM 4.96 3.50 St.Dev JHM 4.02 3.75 JHM $=0$ 5 17 0.007 JHM $\leq 1$ 20 35 0.023 JHM $\leq 2$ 38 51 0.092 JHM $\leq 3$ 50 67 0.028 JHM $\leq 4$ 59 73 0.069 JHM $\leq 5$ 70 80 0.181 $N$ 138 135
@@
We also ask how a subject’s choice consistency, as captured by the above Jaccard-Houtman-Maks index, is related to their response times. Figure 4 (top panel) shows that there is a significant –and very similar in size, with a Spearman coefficient of 0.35 and 0.45– positive correlation in both the forced- and free-choice treatments between subjects’ active-choice Jaccard-Houtman-Maks indices and their average response times. That is, choice consistency is negatively correlated with the time it takes for subjects to make their active choices. This interesting finding is broadly consistent with the important prediction of the drift-diffusion neuro-economic model [^4] See Ratcliff and McKoon (2008), Alós-Ferrer, Fehr, and Netzer (2021) and references therein. whereby shorter response times are more likely when a clearly preferred alternative exists, which would imply in turn more consistent active-choice behaviour. The novelty here in this respect is that such a finding is documented in a rich dataset that comprises both binary and non-binary menus. Another possible explanation for this finding is that subjects with a higher cognitive ability (an unmeasured variable in this study) are both more consistent and faster than subjects with a lower cognitive ability. A distinct possible explanation is that spending more time before deciding is more characteristic of people who are prone to second thoughts and therefore more likely to be involved in choice reversals when presented with a series of decision problems. Conversely, it is also possible that subjects who had neither a stable preference relation over the choice alternatives nor a clear decision rule ended up spending more time on average at each menu, but without this extra time ultimately alleviating the effects of their ambivalence.
@@
5         Individual-Level Analysis
5.1         Non-Parametric Model Fitting
We now turn to the tests of our possibly multi- and/or empty-valued experimental choice data for compatibility with each of the three models introduced in Section 2. These tests were done by applying a model-based extension of the classic Houtman and Maks (1985) that was used in the aggregate-level analysis, and extending it to our generally multi- and/or empty-valued general-choice data. Specifically, we ask what is the smallest number of decisions that need to be dropped/changed in a subject’s dataset in order for it to be compatible with that model under some of its instances/admissible preference orderings. We refer to this number as the model’s distance score. [^5] We use the term “distance score” rather than “distance” partly because the underlying function is not a proper metric and partly to highlight the ranking aspect of this JHM-based goodness-of-fit approach. These scores were computed using a method that was originally introduced in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016, 2022) [^6] The 2016 working-paper version of that study indirectly constructed choice correspondences by augmenting the subjects’ actual choices with their survey responses in indifference-vs-preference questions that followed binary menus only. Before constructing the subjects’ indifference classes from these data, the survey responses were analysed for consistency, and responses that eventually led to intransitive indifferences were discarded. Whenever this was the case, the subjects’ actual choices were interpreted as revealing strict preferences, regardless of the survey response. and which, since Gerasimou and Tejiščák (2018), has been extended considerably and made freely available online in the open-source desktop application [Uncaptioned image]. [^7] With retainment of its original focus on the model of Rational Choice/Utility Maximization with strict preferences, the HM approach has been extended in other intuitive directions by Apesteguia and Ballester (2015) and Dean and Martin (2016). However, although the “Swaps” measure of Apesteguia and Ballester (2015) that pertains to general choice environments is also readily computable in [Uncaptioned image], it is currently unclear how it should be computed when subjects choose multiple items from the same menu, or how it should be extended in models that generalise utility maximization.
More specifically, a brute-force algorithm was implemented for these combinatorial-optimi-zation computations. This involved the production of all choice datasets that are generated by all possible instances of every model, and comparing each such dataset against every subject’s own dataset in order to find the model(s) and instance(s) that are closest to it in the minimum distance-score sense. Therefore, this method detects perfect as well as approximate model fits, and in the latter case it quantifies the approximation in an intuitive way that may be interpreted as the number of “mistakes” made by an agent who might be portrayed as following a particular decision rule. [^8] For other approximations that have been proposed or applied recently for/at distinct classes of models and choice environments we refer the reader to Choi, Fisman, Gale, and Kariv (2007); Echenique, Lee, and Shum (2011); Choi, Kariv, Müller, and Silverman (2014); Apesteguia and Ballester (2015); Dean and Martin (2016); Bouacida and Martin (2021); Dembo, Kariv, Polisson, and Quah (2021); Fudenberg, Liang, and Gao (2023); Cherchye, Demuynck, Lanier, and Rock (2023); Echenique, Imai, and Saito (2023) and references therein.
We stress that despite the vast numbers of weak orders (4,683), partial orders (130,023) and, especially, incomplete preorders (209,527) that are defined on a set of 6 alternatives (OEIS, 2021), the above tool makes such an exact computation possible very quickly –in less than 12 seconds– for all subjects in each treatment, for each of the three models. We also stress that, although this tool lies within the rapidly expanding realm of artificial-intelligence technologies, broadly interpreted, it builds on combinatorial-optimization methods to arrive at exact solutions to a clearly defined and behaviourally intuitive problem. In particular, it does not feature so-called “deep-learning” decision algorithms that build on neural-network structures which may be unclear/uninterpretable or prone to over-fitting. Finally, although the brute-force algorithm is linear in the number of subjects but exponential in the number of alternatives, the model-rich distance-score method itself is scalable and can be extended to analyse datasets that are derived from much larger sets of alternatives by employing powerful open-source constraint solvers.
Undominated Dominant Utility Choice with Choice with Maximization Incomplete Incomplete All Preferences Preferences Forced-Choice treatment ($N=138$) Subjects with score $=$ 0 5 (4%) 0 0 5 (4%) Subjects with score $\leq$ 5 61 (44%) 16 (11.5%) 0 77 (56%) Mean/median best score ($\leq 5$) 2.01/1.83 3.56/4 2.34/2 Mean / median best-model 1.23/1 1.19/1 1.22/1 preference orderings (score $\leq 5$) Minimum score in simulations 14.87 13.92 Non-Forced-Choice treatment ($N=135$) Subjects with score $=$ 0 3 (2%) 0 7 (5%) 10 (7%) Subjects with score $\leq$ 5 30 (22%) 7 (5%) 37 (27%) 74 (55%) Mean/median best score ($\leq 5$) 2.12/1.75 2.37/2.08 1.93/2 2.05/2 Mean/median best-model 1/1 1/1 1.16/1 1.08/1 preference orderings (score $\leq 5$) Minimum score in simulations 21.03 19.67 20.08
(a) Forced-Choice treatment
(b) Free-Choice treatment
For human subjects, Rational Choice often tied with one, but never both, of the other two models. The classification presented in Table 6 broke ties in favour of that model whenever a subject’s JHM distance score was produced by Rational Choice and another model (see the notes of Table 6 for more details). Under this tie-breaking assumption, and in line also with model-approximation methods used in the above-cited studies, the two exhibits show the number, proportion and relative-frequency distributions of distinct subjects that are on average within 10% (equivalently, within 5 decisions) away from being explainable perfectly by an instance of some model, separately for each of the two experimental treatments. We chose this conservative approximation range for two factors: (i) simulations suggest that a JHM distance score of 5 for any of the three models is extremely unlikely to occur randomly in this decision environment (see Table 6 and Figure 7); (ii) for subjects with a JHM score not greater than 5 there is typically only one best-matching preference ordering that explains their behaviour under the respective subject-optimal model.
The total number (proportion) of Forced-Choice subjects with a perfect or 10%-approximate model fit was 5 (4%; all under Rational Choice) and 77 (56%), respectively, with 61 (44%) and 16 (11.5%) of those in the latter group classified under Rational Choice and Undominated Choice with Incomplete Preferences. Furthermore, the total number (%) of Free-Choice subjects with a perfect or 10%-approximate fit was 10 (7%) and 74 (55%), respectively. Thirty (22%), 7 (5%) and 37 (27%) of all subjects in this treatment were categorized under Rational Choice, Undominated and Dominant Choice with Incomplete Preferences, respectively. In addition, 3 and 7 of these subjects were perfectly compatible with the first and third model, respectively, with one subject among the former three and five among the latter seven revealing indifferences in their weakly ordered and weakly preordered preferences.
The model fit is slightly better on average in the Free-Choice treatment than in the Forced-Choice one (mean/median best JHM scores: 2.05/2 vs 2.34/2), although the distributions are not significantly different. Similarly, and in line with the results from the consistency analysis presented in the previous section, the proportion of subjects who were best-matched by one of the two models of consistent active choices (i.e. Rational Choice and Dominant Choice with Incomplete Preferences) is higher in the Free-Choice treatment (49% vs 44%), while the proportion of subjects best-matched by the third model (Undominated Choice with Incomplete Preferences) that predicts generally inconsistent active choices is lower in that treatment (5% vs 11.5%). Moreover, when Rational Choice was not the best-matching model, its distance score was on average 1.4 and 5.7 units higher in the Forced- and Free-Choice treatments, respectively. Finally, the highest JHM distance score of 5 for experimental subjects in this approximate model-fitting analysis is far below the corresponding minimum scores of simulated random-behaving subjects, which ranged between 14 and 21.
Figure 8: All 37 subjects that were best-matched by Dominant Choice with Incomplete Preferences were unlikely to have achieved this classification at random given the menus where they deferred.
Prompted by the greater permissiveness of behaviour in the Free-Choice treatment, moreover, for the 37 subjects who are best-explained by Dominant Choice with Incomplete Preferences we also conducted a new robustness check. By analogy to the way in which the predictive success of Rational Choice on Free-Choice subjects’ active-choice consistency was evaluated in Section 4.4, we generated another set of 10,000 random-behaving simulated subjects for each of the 37 human experimental participants who were best-matched by that model, this time holding fixed in each such subject’s simulations block the menus at which the relevant human participant deferred. This allows for evaluating the predictive success of the Dominant-Choice model’s 10%-approximation range in this decision environment. Figure 8 shows the histogram with JHM distance-score differences for this model between the 2.5th percentile value in each subject-specific 10k-simulations block and the relevant subject’s actual score. For all 37 subjects the difference is positive and typically much greater than 10 score units (i.e. active-choice or deferral decisions). This analysis suggests that Dominant Choice with Incomplete Preferences is not excessively permissive in this environment, and that its good fit is useful towards explaining behaviour and eliciting preferences.
For the vast majority of all subjects that are approximately matched by a model there is a unique weak/strict and complete/incomplete preference ordering generating their minimum JHM distance score under the respective optimal model, with the mean/median number of such orderings being 1.22/1 and 1.08/1 in the Forced- and Free-Choice treatments, respectively. Importantly, for 60 of 77 (78%) and 59 of 74 (80%) subjects in the Forced- and Free-Choice treatments, respectively, the recovered optimal preference relation(s) has –or, in the case of Undominated Choice with Incomplete Preferences, might be interpreted as having– a non-trivial indifference component. This highlights the potential usefulness of eliciting indifferences for revealed preference and welfare analysis.
The multi- and/or empty-valued experimental choice data and the non-parametric optimization goodness-of-fit method that were presented in the previous (sub)sections jointly allow us to test for the potential presence of both these theory-guided distinctions for the first time. Additionally, going beyond the boundaries of the above two theoretical studies which assumed that the analyst observes active choices/deferrals at all menus that can be derived from the grand choice set, this method allows us to carry out this test and recover the individual’s strict preferences and, where relevant, indifferences and incomparabilities from an incomplete collection of menus that comprises only 50 out of the 63 possible ones.
Figure 9: A subject’s complete weak preferences, optimally recovered by the model of Rational Choice.
Figure 10: A subject’s incomplete preferences, optimally recovered by the Undominated-Choice model, with or without indifferences (right and left, respectively).
Figure 11: A subject’s incomplete weak preferences, optimally recovered by the Dominant-Choice model.
We find that a total of 16 subjects (5 in the Free-Choice treatment) out of the 23 who were best-matched by the Undominated-Choice model revealed “regular” incomplete preferences that afford the Eliaz-Ok distinction between indifference and indecisiveness. Figure 10 illustrates one such example, with the graph on the right depicting the revealed incomplete weak preference relation and the graph on the left the observationally equivalent revealed incomplete strict preference relation. We also find that 30 out of the total 37 Free-Choice subjects for whom the Dominant-Choice model provided the best fit (including the 5 subjects for whom this fit was perfect) revealed incomplete preferences with a non-degenerate indifference part. Figure 11 depicts such an example.
Although indifference is generically non-existent for a large class of the benchmark Bewley (2002) class of incomplete preference relations under uncertainty that are defined on a space of bundles of continuous commodities or uncertain acts (Gerasimou, 2018b), this is very much not the case when such preferences are over finitely many discrete options, as in the choice environment analysed in this paper. Indeed, the preceding analysis suggests that out of all 60 (22% of the total) subjects that the two models of incomplete-preference maximization explain optimally in our sample across the two treatments, 36 (63%) exhibited behaviour that allows for recovering a transitive strict preference relation together with a transitive indifference relation and a generally intransitive incomparability/indecisiveness relation.
5.3         Cross-Validation Test of Model-Fitting Analysis
5.4         Satisficing
In addition, we compute and study the average position of each subject’s chosen item(s) in the 50 menus’ list orderings because, intuitively, an unusually low average value of this metric could also be indicative of satisficing behaviour for the subject in question. We find that for an additional 7 and 5 (4.4%) subjects in the Forced- and Free-Choice treatments, respectively, who were not classified as approximately explainable by one of the three deterministic models that were discussed above, the average positions of their chosen item(s) in the 50 menus’ list orderings were strictly below the 2.5% percentile cut-off value of 1.84 that is derived from simulated uniform-random forced choices. These two distinct analyses therefore suggest that a total of 21 subjects (7.7%) who were not model-classified in Table 6 might be thought of as exhibiting a systematic satisficing behaviour.
(a) Forced-Choice treatment
Distribution range: $[1.15,2.25]$
(b) Free-Choice treatment
Distribution range: $[1.74,2.32]$
Note: $p=0.009$; two-sided Mann-Whitney $U$ test.
5.5         Preference for Randomization
5.6         Menu Meta-Visualization and Menu Preferences for Flexibility
The leading theories of menu preferences in the literature portray decision makers as exhibiting either a preference for flexibility (Kreps, 1979; Dekel, Lipman, and Rustichini, 2001) or a preference for commitment (Gul and Pesendorfer, 2001, 2005). [^12] Toussaert (2018) reports on an experimental design that was explicitly constructed to test for the presence of such preferences. The former preference is manifested at a menu $A$ if $B$ is preferred to $A$ whenever $B\supset A$. By contrast, a preference for commitment in the Gul-Pesendorfer theory of temptation and self-control is formalized with the Set Betweeness axiom whereby “$A$ is preferred to $B$” implies “$A$ is preferred to $A\cup B$ is preferred to $B$”.
However, our basic analysis at the aggregate level (see, in particular, Table 4), readily suggests that few -if any- subjects in our sample of eligible subjects are likely to exhibit a systematic preference for flexibility. [^13] We note that a total of 4 subjects in the original sample could indeed be thought of as showing such a consistent preference for flexibility throughout the experiment by always choosing all alternatives at every menu. Those subjects were excluded (see Table 2) from the subsequent analysis because their behaviour was completely uninformative about their preferences over gift-card pairs. This point is validated in Figure 13, which depicts the subjects’ cumulative densities of flexibility violations. More specifically, 5 and 12 are the smallest numbers of menus where this principle was violated, in each case by a single subject, while the means/medians in the two treatments stand at 45.8/48 (Forced-Choice) and 46.6/49 (Non-Forced Choice). Therefore, in line with the caveats expressed at the beginning of this subsection, looking at our experimental data through this lens does not seem to add any notable new insights into subjects’ behaviour. This provides some further reassurance toward the validity of the results presented earlier in this section.
5.7         Related Literature
The core of this paper’s experimental design extends in the multi-valued choice direction the design of Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022). That study also reported on forced- and non-forced choice treatments but only elicited single-valued choices. This made it impossible to raise and answer the novel questions that this paper is mainly concerned with. Although a pre-publication version of that paper from 2016 had proposed a method of multi-valued choice construction, that method is considerably more restrictive than the one advanced here. In particular, contrary to this paper’s unrestricted elicitation of subjects’ multi-valued choices, that method used survey data at binary menus in order to augment with (constructed) transitive indifference relations the subjects’ single-valued choices at binary as well as non-binary menus. In work that followed the above-mentioned study but which predates the present one, Bouacida (2018) proposed a different design that elicits multi-valued choices directly, by paying subjects a fixed extra amount for every additional alternative they choose from a menu of real-effort tasks, assigning them one of these tasks at random as a reward.
The design proposed herein differs from both these pre-existing ones and, in our view, materially so. Specifically, like Bouacida (2018) but unlike Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016), we allow subjects to choose multiple alternatives at every menu directly, thereby avoiding the imposition of a priori consistency constraints on the collected multi-valued choice data. Yet, like Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016) and unlike Bouacida (2018), we penalize inconsistent choices towards incentivizing subjects to reveal their true preferences. Furthermore, Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016) implemented their design on all menus derived from a set of 5 headsets and with a 1-out-of-4 subject reward frequency [^14] A more complete list of the main differences between this experiment and those of that study is the following: (i) It allows for potentially multi-valued choices vs requires necessarily single-valued ones; (ii) Features 6 vs 5 choice alternatives (gift-card bundles vs headphones); (iii) Presents 50 decisions vs 26; (iv) Includes 273 subjects in the sample vs 161 & 121; (v) Gives choice rewards to every subject vs 1 out of 4; (vi) No provision of any information about the options in the randomly selected menu before the 2nd decision vs provision of such information in the 1st experiment of the earlier study., while Bouacida (2018) implemented his on all menus derived from a set of four real-effort tasks where all subjects were rewarded. In addition, there are significant differences in the choice alternatives, number of choice problems, subjects’ reward frequency, and most importantly, in the data-analytic methods, research questions and conclusions of these papers.
Turning to the pre-existing literature on the elicitation of incomplete preferences, this has mainly revolved around choice over binary menus of lotteries and/or uncertain acts, and is reviewed in detail in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022). More summarily here: to identify incompleteness in such environments those studies have used choice-deferral alongside preference-for-flexibility models (Danan and Ziegelmeyer, 2006); partially incentivized methods of imprecise-preference revelation from menu lists (see Cubitt, Navarro-Martinez, and Starmer (2015) and references therein); and incentivized methods where incompleteness can potentially be revealed via certain patterns of randomized choice (Cettolin and Riedl, 2019; Agranov and Ortoleva, 2023). In general choice environments over durable goods on the other hand, strict incomplete preferences were documented in the non-forced-choice treatment of Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022) via costly choice deferrals and restricted application of the model-based Houtman-Maks method to the dominant-choice model with strict incomplete preferences. By allowing for multi-valued choices in both Forced- and Free-Choice treatments, and by proposing and building the analysis upon the novel Jaccard-similarity modified variation of the Houtman-Maks method for multi-valued choice data, the design proposed in the present paper significantly improves upon Costa-Gomes et al. (2022) by enabling the analyst to test the hypothesis of of incomplete-preference maximization in the most general choice environments possible. Applied in these data, and based on the distinct choice-theoretic separations of these relations that were proposed in Eliaz and Ok (2006) and Gerasimou (2018a), the method allowed to document empirically –and for the first time– the uncoupling of revealed indifference and incomparability/incompleteness that is predicted by these models.

### v6

@@ 5.1   Choice Sizes
Turning to patterns in the aggregate data, we start with a key new variable of interest, the subjects’ choice size, which is defined simply as the number of gift-card pairs that were chosen at each menu. This variable ranges between 0 and 4 in the Free-Choice treatment, and between 1 and 4 in the Forced-Choice treatment. However, because choice sizes 1 and 2 (as well as 0 in the former treatment) are always feasible, whereas 3 and 4 are not always so, we adjust their relative frequencies accordingly. Specifically, Figure 6 presents the distributions of menu-size adjusted choice sizes in the two treatments, which are derived once their absolute frequencies are divided by the total number of menus where these choice sizes might be observed. For example, the denominators here are $50\times N$ and $15\times N$ for choice sizes 1 and 4, respectively, where $N$ is the total number of subjects in the relevant treatment.
This comparison shows that the modal adjusted choice size was 1 in both treatments, at a rate of just over 50%, and was followed by sizes 2 and 3. The deferral rate in the Free-Choice treatment, defined as the relative frequency of a zero choice size, is 7.4%. This is slightly higher than the 6.9% choose-everything rate, which is obtained by weighted-averaging the relative frequencies where $n$ out of $n$ items were chosen, for $n=2,3,4$ (Table 6). The latter rate is also lower than the corresponding one in the Forced-Choice treatment (8.5%), and the difference is statistically significant ($p<0.001$; two-sided Fisher’s exact test). This difference is consistent with the intuition that, in the absence of the possibility to delay choice when faced with a potentially difficult decision, subjects are more likely to choose everything that is available in the menu, either as a result of following a more general decision rule or, in the context of our experimental design, possibly due to a preference for randomization. Finally, the rate at which subjects chose all 4 alternatives was very low in both treatments, at 3.2% and 2.5% in the Forced- and Free-Choice treatments, respectively.
@@ 5.2   Response Times
Notably, the average response time in both treatments was always lowest ($\approx$ 6 secs) when subjects chose a single pair of gift cards, while it increased monotonically (and statistically significantly) in similar ways as the size of choices and menus increased (Figure 7). These novel empirical facts might be seen as intuitive evidence suggesting that the subjects’ decision was easier at menus where they chose a single gift-card pair, perhaps because that was their clearly preferred one. The latter explanation would be in line with the one often mentioned in the literature of two-alternative forced-choice decisions via sequential-sampling processes such as the drift-diffusion model and extensions (Ratcliff and McKoon, 2008; Alós-Ferrer et al., 2021), although the novelty of identifying a possibly structure for the evolution of this process in free- and non-binary choices would remain.
@@ 5.3   Choice Proportions
Additionally, 69 subjects (51.1%) in the Free-Choice treatment deferred at least once, with deferrals per subject ranging between 1 and 42 (see absolute subject frequencies in Figure 8-(b); bottom panel), and with a mean/median occurrence of 7.28/5 menus. Moreover, 81 subjects in this treatment (60%) chose all feasible gift-card pairs at least once, with such occurrences per subject ranging between 1 and 38 menus (Figure 8-(b); middle panel), and with a mean/median occurrence of 3.42/1 menus. By contrast, in the Forced-Choice treatment there were 98 subjects (71%) who chose everything at least once, with occurrences per subject ranging between 1 and 45 (Figure 8-(a); middle panel), and with a mean/median of 4.22/2 menus. The difference in the two proportions is borderline (in)significant at the 5% level ($p=0.058$; two-sided Fisher’s exact test). These findings lend further support to the possibility that, unable to delay their choice when faced with a hard decision, subjects are more likely to choose everything.
@@ 5.4   Choice Consistency
For this analysis –which is summarized in Table 7– we compute and compare the JHM scores of subjects’ active choices under Utility Maximization while accounting for the possibility of non-trivial indifferences. That is, we compare subjects’ single- and/or multi-valued choices –ignoring any deferral decisions– that are compatible with indifference-permitting utility maximization. First, we focus on the distributions of the JHM scores across the two treatments. Their mean (5.38 v 4.43), median (4.96 v 3.50) and standard deviation (4.02 v 3.75) are uniformly and significantly ($p=0.045$) lower in the Free-Choice than in the Forced-Choice treatment. Thus, the number of partially or fully “mistaken” –from the point of view of utility maximization– active choices is significantly smaller for subjects who were not asked to always choose some gift-card pair(s). This is also reflected in our second test for treatment effects in active-choice consistency, where we compare the proportions of subjects who made perfectly (JHM score of 0) or approximately (JHM score less than or equal to 1, 2, 3, 4 or 5) consistent active choices in the two treatments. There were approximately 4% and 13% perfectly consistent subjects in the Forced- and Free-Choice treatments, respectively ($p=0.007$). The pattern is similar for approximately consistent subjects, with 14.5% v 26% making up to one fully mistaken choice ($p=0.023$), 27.5% v 37% up to two ($p=0.092$), 36% v 50% up to three ($p=0.028$), 43% v 54% up to four ($p=0.069$) and 51% v 59% up to five ($p=0.181$).
@@
Forced Choice Free Choice $p$-value Mean JHM 5.38 4.43 0.045 Median JHM 4.96 3.50 St.Dev JHM 4.02 3.75 JHM =0=0 5 17 0.007 JHM $\leq 1$ 20 35 0.023 JHM $\leq 2$ 38 51 0.092 JHM $\leq 3$ 50 67 0.028 JHM $\leq 4$ 59 73 0.069 JHM $\leq 5$ 70 80 0.181 $N$ 138 135
@@
We also ask how a subject’s choice consistency, as captured by the above Jaccard-Houtman-Maks index, is related to their response times. Figure 9 (top panel) shows that there is a significant –and very similar in size, with a Spearman coefficient of 0.35 and 0.45– positive correlation in both the forced- and free-choice treatments between subjects’ active-choice Jaccard-Houtman-Maks indices and their average response times. That is, choice consistency is negatively correlated with the time it takes for subjects to make their active choices. This interesting finding is broadly consistent with the important prediction of the drift-diffusion model mentioned earlier, which predicts that shorter response times are more likely when a clearly preferred alternative exists, which would imply in turn more consistent active-choice behaviour. The novelty here in this respect is that such a finding is documented in a rich dataset that comprises both binary and non-binary menus. Another possible explanation for this finding is that subjects with a higher cognitive ability (an unmeasured variable in this study) are both more consistent and faster than subjects with a lower cognitive ability. A distinct possible explanation is that spending more time before deciding is more characteristic of people who are prone to second thoughts and therefore more likely to be involved in choice reversals when presented with a series of decision problems. Conversely, it is also possible that subjects who had neither a stable preference relation over the choice alternatives nor a clear decision rule ended up spending more time on average at each menu, but without this extra time ultimately alleviating the effects of their ambivalence.